OpenStudy (anonymous):

1/2x+1/4x=16 solve for X

OpenStudy (anonymous):



OpenStudy (anonymous):

Someone just asked this a minute ago, and I answered it for them: (1/2)x + (1/4)x = 16 2x + x = 64 3x = 64 x = 64/3 or 21 1/3 Answer: 64/3 or 21 1/3

OpenStudy (anonymous):

The answer is 21.3

OpenStudy (austinl):

\(\dfrac{1}{2}x+\dfrac{1}{4}x=16\) \(\dfrac{2x}{4}+\dfrac{x}{4}=16~\Rightarrow~\dfrac{3x}{4}=16\) Divide both sides by \(\dfrac{3}{4}\) which is the same as multiplying by \(\dfrac{4}{3}\). \(x=16\times\dfrac{4}{3}=\dfrac{64}{3}\)

OpenStudy (anonymous):

if x is in the neumerator .. @greenlegodude57 is rt i took x in the denomenator .. @wolowizzard

OpenStudy (anonymous):

My answer is also correct, just as yours is. There are 2 methods of doing this. I did it one way you did it the other. We are both correct.

OpenStudy (austinl):

@greenlegodude57 The answer is not 21.3 That is simply an approximation. It is either \(\dfrac{64}{3}\) OR \(21\dfrac{1}{3}\). \(21.\overline{3}\) is not exact.

OpenStudy (anonymous):

Yes, but it's a choice, in his multiple choice question, I told you I did this before a few minutes ago.

OpenStudy (anonymous):

Look: http://openstudy.com/study#/updates/5242d741e4b0838bd448a1fa

OpenStudy (anonymous):

21.3 is the CLOSEST answer even though it is not exact.

OpenStudy (anonymous):

(That is a link to the exact problem but with the choices)

OpenStudy (austinl):

I will say this one last thing. The asker did not specify that it was multiple choice, and we should not assume that it was. We were given a problem, that as is, should be solved in exact terms. You should always encourage the use of exact terms, because that is what is expected later on in mathematics. ~Austin out

OpenStudy (anonymous):

I know the asker didn't specify that it was a multiple choice, but once again I told you I did that question a few minutes ago as I put in a link.